Hello,

I need help finding the surface area of a cylinder with a 10m radius and a 12m height.
Thank you

Answer by ganit(6)

surface area of a cylinder with a 10m radius and a 12m height.
The formula for surface area = 2 × r × pi × L + 2 × pi × r^2.
where r= radius and L= height

Here r = 10 m and L= 12 m , pi =3.14

Substituting the values in the formula
Surface area = 2* 10 *3.14 + 2*3.14*10 *10
= 62.8 + 628
= 690.8 sq m
Surface area of the cylinder = 690.8 sq m